OpenStudy (anonymous):

find the solution to the quadratic equation by completing the square x^2+6x=7 @siegel

OpenStudy (anonymous):

well first get the squared part alone, so how do you do that?

OpenStudy (anonymous):

um subract 6x from both sides?

OpenStudy (anonymous):

right so what is your equation now?

OpenStudy (anonymous):

x^2=7-6x

OpenStudy (anonymous):

\[x^{2} + 6x - 7 = 0\] \[(x^{2}+6x-9+9) - 7 = 0\] \[(x+3)^{2}-9-7 = 0\] \[(x+3)^{2}-16 = 0\] Therefore x = -3 and y = -16 (-3,-16) is the solution!

OpenStudy (anonymous):

wait I did something wrong, the other person got it, would you like me to walk you through it?

OpenStudy (anonymous):

yeah because i still dont understand it...

OpenStudy (anonymous):

Okay, sorry about my slip up. keep the equation the same but think about FOIL. we are focusing on the left side, you have an a value and a b value, how would you make it a quadratic equation?

OpenStudy (anonymous):

the equation would have to =0? so then i guess you would have to subtract 7 from both sides.

OpenStudy (anonymous):

not exactly, remember we are only dealing with the left side, how would you make\[X ^{2}+6X\] into a perfect square trinomial, you can add anything. wait do you know about perfect square trinomeals?

OpenStudy (anonymous):

9?

OpenStudy (anonymous):

right, so how did you get that?

OpenStudy (anonymous):

well if you factored it using FOIL and knowing it was a perfect square the last term would have to be a perfect square...and the only number that adds up to 6 being a perfect square is 3. if that makes any sence

OpenStudy (anonymous):

No you are totaly on track. a quicker way is to divide the b cooeficient by 2 then square it. now what does your equation look like?

OpenStudy (anonymous):

\[x ^{2}+6x+9=7\]

OpenStudy (anonymous):

close but remember what you do to one side you have to do to the other

OpenStudy (anonymous):

so wait then you would multiply it by 3? making it 21?

OpenStudy (anonymous):

no you added 9 to one side to make it a pst, so wouldn't you add 9 to the other side?

OpenStudy (anonymous):

oh! now i understand! so then the equation would be x^2+6x+9=16

OpenStudy (anonymous):

right! Okay, so remember how you had that 3, well what is the factor of x^2+6x+9

OpenStudy (anonymous):

ok..

OpenStudy (anonymous):

factors sorry.

OpenStudy (anonymous):

okay so then what do you do?

OpenStudy (anonymous):

well does the equation look like now?

OpenStudy (anonymous):

x^2+6x+9=16

OpenStudy (anonymous):

now factor the left side

OpenStudy (anonymous):

I promise this will all make sense.

OpenStudy (anonymous):

(x+3)^2

OpenStudy (anonymous):

=16

OpenStudy (anonymous):

so plug that into the equation

OpenStudy (anonymous):

into the quadratic equation?

OpenStudy (anonymous):

well if (X+3)^2=16 how could we get rid of the ^2?

OpenStudy (anonymous):

by.....taking the square root from both sides?

OpenStudy (anonymous):

right! so what do we have now?

OpenStudy (anonymous):

x+3=4

OpenStudy (anonymous):

yes the only thing is, what do you do when you take the square root of a number, (this will give you the two answers)

OpenStudy (anonymous):

you have a positive and a negative answer

OpenStudy (anonymous):

yes so what is the equation?

OpenStudy (anonymous):

well you would have the same equation but when you solve it you would end up with 1 and negative 7

OpenStudy (anonymous):

shoot class is over. I'll be back though if you are still on. thanks :)

OpenStudy (anonymous):

YES! after taking the square root of both sides you have\[(X+3)=\pm4\] +4-3=1 and -4-3=-7. You are awesome! great job

OpenStudy (anonymous):

awesome!!! thanks :)

OpenStudy (anonymous):

No problem any time!

OpenStudy (anonymous):

okay same concept but different equations. can you help? x^2+12x=-11

OpenStudy (anonymous):

ok so how do you make the left side a pst

OpenStudy (anonymous):

6?

OpenStudy (anonymous):

oh wait I think i figured it out...the answer would be x=-1 and-11?

OpenStudy (anonymous):

what about the ones like this? x^2+2x-4=0?

OpenStudy (anonymous):

wait the last one how do you get -1 and -11, I think you have to switch the signs. for the next one how do you make the left side a binomial?

OpenStudy (anonymous):

subtract 4 from both sides?

OpenStudy (anonymous):

close it's a negative 4 so how do you eliminate it?

OpenStudy (anonymous):

you would subtract -4 from both sides so your equation would be...x^2+2x=-4

OpenStudy (anonymous):

I see what you're thinking but you're subtracting -4 from both sides, what do you do when you subtract a negative?

OpenStudy (anonymous):

im not sure...switch the sign?

OpenStudy (anonymous):

5-(-7) = 5+7

OpenStudy (anonymous):

oh they cancel!

OpenStudy (anonymous):

yep so what does the equation look like now?

OpenStudy (anonymous):

x^2+2x=4

OpenStudy (anonymous):

right! so what comes next?

OpenStudy (anonymous):

you would have to turn it into a perfect square. so the equation would be..x^2+2x+4=8

OpenStudy (anonymous):

close, what do you do to the b value?

OpenStudy (anonymous):

wait you dont do anything to B.

OpenStudy (anonymous):

go back to the binomial. x^2+2x=4 divide the b value by 2 and then square it.

OpenStudy (anonymous):

that would be 1

OpenStudy (anonymous):

right so what is the equation

OpenStudy (anonymous):

x^2+x+4=8?

OpenStudy (anonymous):

no...remember the number you get has to make a pst. 1 will end up being the c value. try factering it with 1 as the a value, 2 as the b value and 1 as the c value.

OpenStudy (anonymous):

hahaha wait im totally confused.

OpenStudy (anonymous):

ok, so in this problem the answer will not be perfect, but go with it. start from the beginning, you add 4 to both sides x^2+2x=4 now how do you make (x^2+2x) a perfect square trinomial

OpenStudy (anonymous):

by adding 4 to both sides

OpenStudy (anonymous):

no, although I agree, that is the logical first choice. but remember, the c value times 2 has to be equal to the b value

OpenStudy (anonymous):

so if it has to be the same then wouldnt it be 2?

OpenStudy (anonymous):

b is always twice the square root of c so c is ^2 of half of b

OpenStudy (anonymous):

hmm......that would be 1

OpenStudy (anonymous):

yes, this is a bad example but yes. so write out the equation

OpenStudy (anonymous):

x^2+2x+1=4

OpenStudy (anonymous):

check the right side

OpenStudy (anonymous):

=5

OpenStudy (anonymous):

right, so what's next?

OpenStudy (anonymous):

put in quadratic?

OpenStudy (anonymous):

well factor the left side

OpenStudy (anonymous):

oh so then it would be x+1)=5

OpenStudy (anonymous):

Squared

OpenStudy (anonymous):

like (x+1)^2=5

OpenStudy (anonymous):

great what do you do next

OpenStudy (anonymous):

take the square root from both sides

OpenStudy (anonymous):

so what happens to the right side PS keep it in radical form.

OpenStudy (anonymous):

it would be x+1=positive and negative \[\sqrt{5}\]

OpenStudy (anonymous):

sorry if im like bugging you but I have two more problems that are the same type that Im not sure about... 2x^2+3x-5=0 I cant make this a perfect square

OpenStudy (anonymous):

No problem also with x^2+12x=-11 I think you need to switch your signs for your answer. ok this is something that gets easier with practice. what do you think you do first

OpenStudy (anonymous):

you have to get the parts with the x alone. so you would add 5 to both sides making the equation 2x^2+3x=5

OpenStudy (anonymous):

actually for this one the first thing you do is get the cooeficient of 2x^2 to 1. how do you think you do that? technically you can move the five first just remember to do the same thing to both sides.

OpenStudy (anonymous):

what do you mean get it to 1?

OpenStudy (anonymous):

you are going to divide everything by 2. I know this is weird but it's easier with even numbers.

OpenStudy (anonymous):

but that seems weird because then you have x^2+3x

OpenStudy (anonymous):

well you divide EVERYTHING by 2 so 3 becomes 3/2 or 1.5

OpenStudy (anonymous):

so then you have x^2+1.5x=2.5?

OpenStudy (anonymous):

exactly, here you're going to want a calculator. I don't know if your teacher want's it in decimal or fraction form so use your best judgement. What do you think you do next?